# Cascade-driven opinion dynamics on social networks

### v2

@@ III-D Polarization
In our previous work [7], we reviewed and classified the most popular polarization metrics used in the related literature. These metrics, given an opinion vector, measure polarization according to different semantics. They can be summarized in the following definition.
@@
$P_{4}(\bm{x})$ $=\sum_{i\in\mathcal{V}}|x_{i}|=\|\bm{x}\|_{1}$ (13)
In the context of measuring polarization in social networks, the metrics $NDI$, GDIGDI, and $P_{1}$ quantify polarization based on the degree of disagreement among individuals’ opinions. On the other hand, metrics $P_{2}$, $P_{3}$, and $P_{4}$ consider polarization to occur when opinions are significantly distant from a neutral opinion, which is assigned a value of 0.
@@
is referred to as the polarization value and, by definition, is positive if and only if the model is polarizing with respect to the metric $\Phi$. Our previous research [7] showed that the FJ model is inherently polarizing and provided formulas to derive polarizing vectors for the polarization metrics $P_{2}$, $P_{3}$, and $P_{4}$. These vectors are closely linked to the spectral properties of the matrix $H=(I-\Lambda W)^{-1}(I-\Lambda)$, which maps initial opinions $\bm{u}$ to final opinions $\bm{z}$ as in Equation (2). Depending on the considered sub-domain of the opinion space, different polarizing vectors can be identified. In this paper, for $P_{4}$ polarization we use $\bm{u}_{B_{1}(1)}$, while for $P_{2}$ and $P_{3}$ polarization we rely on $\bm{u}_{B_{2}(1)}$, $\bm{u}_{B_{2}(t)}$, and the heuristic $\bm{u}_{V_{>1}}^{heu}$, listed in order of increasing polarization strength. For a detailed mathematical treatment and derivations, we refer the reader to [7]. In Section V we will examine the impact of information cascades when the opinion dynamic process starts with users featuring these polarizing opinions. Notably, we will observe that FJC model induces significantly greater polarization than the FJ for most of the scenarios studied.
@@ IV The Friedkin-Johnsen on Cascades model
In this section, we will discuss the FJC (Friedkin-Johnsen on Cascades) model. The FJ model is an opinion dynamic model that explains how people form their opinions in the offline world, where face-to-face interactions significantly influence the depth of relationships and consequently the shared understanding of opinions. This is the unique mechanism that triggers the tension resolving process, which is the brain’s reaction to cognitive dissonance. Information cascades, on the other hand, describe how information is spread through online social networks, where interactions are encoded as engagement with each other’s posts. They characterise how a single post by a source can go viral and inundate the entire network. Despite belonging to two different and separate research areas, opinion dynamics and information cascades are intricately linked. In online social networks, opinions are expressed in a cascade manner, by subsequent exposure to other contents/opinions. It is reasonable to expect that this unique exposure to others’ posts influences how we update our opinions. The amplification of virality in social networks can disrupt the balance of real-life discussions in offline networks, thus strengthening the role of central nodes. This is a crucial insight that has far-reaching implications for understanding opinion formation and information dissemination in online social networks.
IV-A Revisiting the independent cascade model
The standard independent cascade model needs to be slightly adjusted to integrate it into the FJC framework. To this aim, we identify three new node states, which replace the previous ones: unaware, aware and spreader. A node is unaware when it has not been exposed to the post yet. As soon as it “sees” the post from a neighbour, the node becomes aware and then it updates its opinion as described in the following sections. With a probability $\theta_{s,i}$ it also reshares the post, and, when it does, the node moves into the spreader state. An illustration of the possible transition of states for a single node is provided in Figure 1. We assume (A1) that the content being shared does not express an opinion on its own, but rather the opinion is conveyed only by the person sharing it along with a comment. This assumption allows us to avoid considering the post as an external factor that can modify opinions and instead to focus on studying how the simple cascade dynamic affects opinion formation in comparison to the FJ scenario.
@@ IV-A Revisiting the independent cascade model
In our model, an individual identified as the source disseminates a piece of news to the online social network. In the rest of this paper, we streamline our discussion by initially constructing the model under the assumption that each post originates from a single seed node. However, it is worth noting that the model’s flexibility allows for straightforward extension to accommodate multiple sources, akin to the original information cascade model. In fact, including multiple sources simply affects the first level of the information cascade tree, while the rest remains unchanged. We assume (A2) that once the post is shared the first time by the seed node $r$, it spreads outwards like a tree, from the original source to the first, second, third, and subsequent in-neighbours, which we will refer to as the nodes in the first, second, third, … layer. This assumption means that the opinions of parent nodes represent the first exposure a node has in its news feed, thereby solidifying their influence. Following this initial encounter, we maintain that nodes typically become disinterested in posts shared by others. Thus, we refer to the $l$-th in-neighbourhood as the $l$-th layer (or layer $l$), and denote it by $\mathcal{O}_{l}^{(r)}$. The depth of the graph is given by the eccentricity of the root node, denoted by $\epsilon^{(r)}$, which corresponds to the maximum distance from $r$ to any other node. Accordingly, the entire tree comprising all nodes in the aware state of the cascade is represented as the sequence $\mathcal{O}^{(r)}=\left(\mathcal{O}l^{(r)}\right)_{l=0}^{\epsilon^{(r)}}$ of all layers.
@@
Figure 2: A toy example of a network in (a) and the corresponding propagation tree for $r=1$ in (b). Nodes 22 and $3$ update their opinion deterministically, while the others, do it with probability $\theta$.
@@
Please, remember that $\mathcal{N}^{-}(i)$ refers to the in-neighbours of node $i$.
@@ IV-B The cascade-driven Friedkin-Johnsen
Now, we want to discuss the details of the opinion updating operation. Please observe that, since we want to update opinion layer by layer, we need an asynchronous version of the FJ, similar to that of Equation (3). We focus on the spreading process of the $n$-th post $p_{n}$ generated at node $r_{n}$ which involves nodes in the cascade $\mathcal{C}^{(r_{n})}$. As the cascade starts, the opinions of nodes are equal to $\bm{u}=\bm{x}(0)$ if $n=1$, i.e. if we are considering the spreading of the first post. Otherwise, they are equal to the final opinion of nodes as the previous post $p_{n-1}$ completed spreading. Every time a node $j$ in the cascade reshares the post $p_{n}$ at a time $t$, it conveys its current opinion $x_{j}(t)$ together with the reposting. A node $i$ in layer $l$ can receive the post (which, as explained above, is a proxy for a node’s opinion) from one or more predecessors $j$. This triggers the update of node $i$’s opinion, which is carried out according to the following update rule:
@@
$=\frac{1}{|\varphi(i)|}\Bigg{\{}\sum_{j\in\varphi(i)}h_{i}\left[(1-\gamma_{ij})x_{i}(t-1)+\right$.
$\qquad\left.+\gamma_{ij}x_{j}(t-1)\right]+(1-h_{i})u_{i}\Bigg{\}},$ (17)
where $h_{i}$ and $\gamma_{i}$ are defined as in Equations 4 and 5, and $\varphi(i)$ is defined in the Equation (16). The above update rule factors in the average total influence on node $i$ by all its in-neighbours that have reshared the post in the previous layer. This means that the average is calculated only from those neighbors who have already engaged with the post and shared their opinions, thereby influencing the opinion of node $i$. Note that, differently from the asynchrounous FJ update in Equation (3), node $i$ assesses the opinions of its predecessors all at once, not in a pairwise fashion.
@@ V-A Common experimental settings
All networks used are unweighted, so the social matrix $\hat{W}$ corresponds to the adjacency matrix. As described in Section III-A, the influence matrix is derived by row-normalizing $\hat{W}$: $w_{ij}=\frac{\hat{w}{ij}}{\sum_{k}\hat{w}{ik}}$. Following our previous work [7], we considered three strategies for setting node susceptibility. The first two rely on PageRank centrality, a measure that effectively captures node influence [25]. For each node $i$ with PageRank score $\mathrm{P}_{i}$, we assigned $\lambda_{i}$ either the rescaled value of $\mathrm{P}_{i}$ or $\mathrm{P}_{i}^{-1}$, both mapped to the interval $(0,1)$, thus giving higher susceptibility to more (or less) central nodes. In the third strategy, all nodes were assigned a fixed susceptibility value of 0.6. These variations allow us to explore how susceptibility interacts with node centrality. Finally, since our primary goal is to study the impact of cascades on polarization, we initialized opinions using the polarizing vectors defined in Section III-D, ensuring a consistent and controlled setup for measuring polarization dynamics. Final opinions under cascades were then compared with those of the FJ model, whose outcomes are analytically derived via Equation (2) using the same parameters, except for the reposting probability $\theta$, which is not applicable to FJ.
@@ V-B1 Dataset description and experimental setting
We validate the model using a real-world dataset from Bhowmick et al.  [46], which contains Twitter (now X) data related to the 2015 Nepal earthquake. This dataset is well-suited for our analysis as it includes both the follower network and tweet/retweet histories. Unlike the synthetic networks in Sec. V-C, this is a directed graph, with multiple connected components. Since FJ-based models operate within a connected structure, we extract the largest strongly connected component (LSCC). Some cascades span disconnected components, suggesting hidden links, so we restrict our analysis to cascades occurring entirely within the LSCC. The resulting graph contains 220,974 nodes, 16,772,693 edges, and 32,313 cascades. Initial opinions are set using the polarizing vectors described in Section III-D. However, their computation requires to compute of the eigensystem of an $n\times n$ dense matrix, which becomes impractical due to the extensive size of the network, even when employing optimized algorithms. To tackle this challenge, we derived a subgraph known as Nepal top cascades, which includes nodes from the 500 longest cascading events. The rationale behind focusing on the longest cascades is that they inherently present a greater challenge for our model, as they involve more complex and prolonged interactions among nodes and increase the likelihood that cascades can overlap (thus, violating assumption (A3)). Specifically, we ensured that all nodes participating in these cascades are included in the subgraph using a heuristic for the directed Steiner Tree problem [47]. Table I summarizes the key properties of this subgraph, which encompasses over 6,000 cascades.
TABLE I: Summary of the real-world Nepal dataset subgraphs used for validation
@@
Our goal is to evaluate how well this real dynamic is captured by the proposed FJC model (we also include the classical FJ model for reference). For the FJC model, we need to estimate the resharing probability $\theta_{i}$ for each node. Inspired by weighted cascade models [36, 48], we define $\theta_{i}$ as the ratio between the number of posts reshared by node $i$ and the number of posts it has seen from its followees. Table II shows the distribution of $\theta_{i}$, which is generally low, indicating that most nodes rarely reshare posts. Please note that these values are consistent with the retweet probabilities in [49], remaining below 0.02. Unlike the real dynamics, the FJC model processes cascades in a tree manner and sequentially. Specifically, posts are propagated from the sources to the network layer-by-layer, in accordance with assumption (A1) and each cascade starts only after the previous one ends, in accordance with assumption (A3).
TABLE II: Statistics of followers, followed and $\theta$ in the Nepal subgraph
@@ V-B2 Opinion evolution of real opinion dynamics vs FJC and FJ
Due to the large size of the networks, we cannot visualize the polarizing vectors, since the plots would be unreadable. Instead, our main objective with these datasets is to evaluate the robustness of the FJC model against the real opinion dynamics. We do so by comparing the distributions of opinion shifts—i.e., the differences between final and initial opinions—for the real evolution (labeled Real) and the predictions of FJC and FJ. As shown in Figure 3, FJC closely mirrors the real dynamics observed in the Nepal top cascades dataset, outperforming the standard FJ model. Please note that long cascades are more likely to violate assumption (A3), reducing FJC’s accuracy. Nevertheless, FJC remains surprisingly reliable even in this more challenging scenario.

### v3

@@ III-D Polarization
In [11], the most widely used polarization metrics in the related literature are discussed and classified according to their underlying semantics. We summarize below a selected subset of these definitions, which are used in Section V.
@@
$P_{4}(\bm{x})$ $=\sum_{i\in\mathcal{V}}|x_{i}|=\|\bm{x}\|_{1}.$ (10)
In the context of measuring polarization in social networks, these metrics consider polarization to occur when opinions are significantly distant from a neutral opinion, which is assigned a value of 0.
@@
is referred to as the polarization value and, by definition, is positive if and only if the model is polarizing with respect to the metric $\Phi$. Our previous research [11] showed that the FJ model is inherently polarizing and provided formulas to derive polarizing vectors for the polarization metrics descibed above. These vectors are closely linked to the spectral properties of the matrix $H=(I-\Lambda W)^{-1}(I-\Lambda)$, which maps initial opinions $\bm{u}$ to final opinions $\bm{z}$ as in Equation (2). In this paper, we use the following polarizing initial conditions:
  • $\bm{u}_{B_{1}(1)}$: maximizes polarization under $P_{4}$ within the $L_{1}$-ball of radius 1;
  • $\bm{u}_{B_{2}(1)}$ and $\bm{u}_{B_{2}(t)}$: maximize polarization under $P_{2}$ and $P_{3}$ within the $L_{2}$-balls of radius 1 and $t$, respectively;
  • $\bm{u}_{V_{>1}}^{heu}$: a heuristic approximation that maximizes $P_{2}$ and $P_{3}$ over a larger subdomain of the opinion space.
Analysis of these vectors in [11] revealed which nodes contribute most to polarization shifts in the network. Polarization is maximized when extreme opinions are assigned to the most influential individuals in the network, while less influential or more susceptible individuals are initialized with neutral opinions. Accordingly, these vectors place opinions close to the boundaries of the opinion domain (i.e., near $\pm 1$) on nodes that are either highly PageRank-central or highly stubborn, while assigning opinions close to zero to nodes that are more susceptible and/or less PageRank-central. The analysis in [11] confirmed that PageRank centrality is the centrality measure that better captures the influence among nodes in social networks [43].
In Section V, we focus on these polarizing initial conditions to study the effect of information cascades under worst-case polarization scenarios for the baseline FJ model. While alternative initial opinion profiles would generally lead to different quantitative outcomes, these configurations are known to maximize polarization under FJ dynamics. While they are not guaranteed to play the same role under the FJC dynamics, they provide a principled reference point to assess how the introduction of information cascades modifies polarization. As we will show, information cascades tend to amplify the influence of already central individuals, who are precisely emphasized by the polarizing configurations considered here. For this reason, these scenarios provide a principled and informative reference point to assess how cascades can amplify or mitigate polarization.
@@ IV The Friedkin-Johnsen on Cascades model
In this section, we introduce the Friedkin–Johnsen on Cascades (FJC) model, which combines opinion updating with cascade-based information diffusion. The FJ model describes opinion formation through interpersonal influence, while information cascades describe the propagation of content in online networks. The FJC model explicitly integrates these mechanisms, capturing how opinions are updated in response to exposure to posts in a cascade. This framework provides a basis for studying how information diffusion interacts with opinion dynamics in networked environments. Throughout this section, we introduce several modeling assumptions, which are discussed where they become relevant in order to facilitate understanding of the proposed framework. For the reader’s convenience, all assumptions are summarized in Table I.
TABLE I: Summary of modeling assumptions in the FJC framework
ID Assumption Description
A1 Opinion-neutral content Shared content does not carry an intrinsic opinion; opinions are conveyed only by the individuals sharing it.
A2 Tree-structured exposure A post propagates from the seed in a tree-like manner across successive neighbour layers, with each node influenced only upon first exposure.
A3 Sequential cascade propagation Only one post propagates at a time; simultaneous cascades are not considered.
IV-A Integrating Independent Cascades into the Friedkin–Johnsen Framework
The standard IC model is adapted here to integrate opinion updating into the cascade process, yielding what we refer to as the FJC framework. Rather than restating the full IC mechanism, we explicitly highlight the key differences with respect to the classical model. First, nodes transition through three states—unaware, aware, and spreader. Becoming aware of a post triggers an opinion update, a feature absent in the standard IC model (see Figure 1). Second, information propagation is constrained to a directed tree structure rooted at the source of the post. Nodes are organized into successive neighbourhoods of increasing distance from the source, which we refer to as layers. Under this assumption, each node updates its opinion only upon its first exposure to the post, that is, when it is reached by resharing neighbours in the immediately preceding layer. During this exposure event, a node may aggregate influence from all such neighbours, while subsequent resharings do not induce further opinion updates. As a consequence, the cascade unfolds over a directed tree in which each node appears in at most one layer and is influenced only by its predecessors.
We assume (A1) that the content being shared does not express an opinion on its own, but rather the opinion is conveyed only by the person sharing it along with a comment. This assumption allows us to avoid considering the post as an external factor that can modify opinions and instead to focus on studying how the simple cascade dynamic affects opinion formation in comparison to the FJ scenario.
@@ IV-A Integrating Independent Cascades into the Friedkin–Johnsen Framework
In our model, an individual identified as the source disseminates a piece of news to the online social network. In the rest of this paper, we streamline our discussion by initially constructing the model under the assumption that each post originates from a single seed node. However, it is worth noting that the model’s flexibility allows for straightforward extension to accommodate multiple sources, akin to the original information cascade model. In fact, including multiple sources simply affects the first level of the information cascade tree, while the rest remains unchanged. We assume (A2) that once the post is shared the first time by the seed node $r$, it spreads outwards like a tree, from the original source to the first, second, third, and subsequent follower neighbours (i.e. in-neigbours according to notation of Section III-A), which we will refer to as the nodes in the first, second, third, … layer. Under this assumption, each node can appear in at most one layer of the cascade tree, corresponding to its first exposure to the post. While a node may be influenced by multiple predecessors in the previous layer, all such influences occur simultaneously at this first exposure. Following this initial encounter, we maintain that nodes typically become disinterested in posts shared by others. Thus, we refer to the $l$-th in-neighbourhood as the $l$-th layer (or layer $l$), and denote it by $\mathcal{O}_{l}^{(r)}$. The depth of the graph is given by the eccentricity of the root node, denoted by $\epsilon^{(r)}$, which corresponds to the maximum distance from $r$ to any other node. Accordingly, the entire tree comprising all nodes in the aware state of the cascade is represented as the sequence $\mathcal{O}^{(r)}=\left(\mathcal{O}l^{(r)}\right)_{l=0}^{\epsilon^{(r)}}$ of all layers. Please note that, unlike the standard IC model, where influence attempts may occur independently along multiple incoming edges, here the cascade is explicitly represented as a directed tree rooted at the source, ensuring a unique first exposure for each node.
@@
Figure 2: A toy example of a network in (a) and the corresponding propagation tree for $r=1$ in (b). Nodes 22 and $3$ update their opinion deterministically, while the others do it with probability $\theta$.
@@
Please, remember that $\mathcal{N}^{-}(i)$ refers to the in-neighbours of node $i$. This construction intentionally departs from the classical IC formulation to isolate the effect of single-exposure, cascade-driven opinion updates, which is essential for coupling the cascade process with the Friedkin-Johnsen dynamics.
@@ IV-B The Cascade-Driven Friedkin-Johnsen
Now, we want to discuss the details of the opinion updating operation. Please observe that, since we want to update opinion layer by layer, we need an asynchronous version of the FJ, similar to that of Equation (3). We focus on the spreading process of the $n$-th post $p_{n}$ generated at node $r_{n}$, which involves nodes in the cascade $\mathcal{C}^{(r_{n})}$. As the cascade starts, the opinions of nodes are equal to $\bm{u}=\bm{x}(0)$ if $n=1$, i.e. if we are considering the spreading of the first post. Otherwise, they are equal to the final opinion of nodes as the previous post $p_{n-1}$ completed spreading. Every time a node $j$ in the cascade reshares the post $p_{n}$ at a time $t$, it conveys its current opinion $x_{j}(t)$ together with the reposting. A node $i$ in layer $l$ can receive the post (which, as explained above, is a proxy for a node’s opinion) from one or more predecessors $j$. This triggers the update of node $i$’s opinion, which is carried out according to the following update rule:
@@
$=\frac{1}{|\varphi(i)|}\Bigg\{\sum_{j\in\varphi(i)}h_{i}\Big[(1-\gamma_{ij})x_{i}(t-1)+\gamma_{ij}x_{j}(t-1)\Big]+$
$\hskip 18.49988pt\hskip 18.49988pt+(1-h_{i})u_{i}\Bigg\}\vskip-5.0pt$ (14)
where $h_{i}$ and $\gamma_{i}$ are defined as in Equations (4) and (7), and $\varphi(i)$ is defined in the Equation (13). The above update rule factors in the average total influence on node $i$ by all its in-neighbours that have reshared the post in the previous layer. This means that the average is calculated only from those neighbors who have already engaged with the post and shared their opinions, thereby influencing the opinion of node $i$. Note that, differently from the asynchronous FJ update in Equation (3), node $i$ assesses the opinions of its predecessors all at once, not in a pairwise fashion.
@@ V-A Common Experimental Settings
All networks used are unweighted, so the social matrix $\hat{W}$ corresponds to the adjacency matrix. As described in Section III-A, the influence matrix is derived by row-normalizing $\hat{W}$: $w_{ij}=\frac{\hat{w}{ij}}{\sum_{k}\hat{w}{ik}}$. Following our previous work [11], we considered three strategies for assigning node susceptibility. In the first two, susceptibility is coupled to PageRank centrality, which is known to be strongly correlated with node influence in linear opinion dynamics, including the Friedkin–Johnsen model [43]. While PageRank does not exactly coincide with the influence weights induced by FJ dynamics, it provides a well-established proxy for identifying influential nodes in directed networks. Specifically, for each node $i$ with PageRank score $\mathrm{P}_{i}$, we assign susceptibility $\lambda_{i}$ either proportional to $\mathrm{P}_{i}$ or to $\mathrm{P}_{i}^{-1}$, with values rescaled to the interval $(0,1)$. These two configurations allow us to study stressed scenarios in which susceptibility is systematically aligned or anti-aligned with node influence. To assess the role of this coupling, we also consider a control setting in which all nodes are assigned the same susceptibility value, $\lambda_{i}=0.6$, thereby fully decoupling susceptibility from centrality. Finally, since our primary goal is to study the impact of cascades on polarization, we initialized opinions using the polarizing vectors defined in Section III-D, ensuring a consistent and controlled setup for measuring polarization dynamics. Final opinions under cascades were then compared with those of the FJ model, whose outcomes are analytically derived via Equation (2) using the same parameters, except for the reposting probability $\theta$, which is not applicable to FJ.
@@ V-B1 Dataset description and experimental setting
We validate the model using a real-world dataset from Bhowmick et al.  [8], which contains Twitter (now X) data related to the 2015 Nepal earthquake. This dataset is well-suited for our analysis as it includes both the follower network and tweet/retweet histories. Unlike the synthetic networks in Sec. V-C, this is a directed graph, with multiple connected components. Since FJ-based models operate within a connected structure, we extract the largest strongly connected component (LSCC). Some cascades span disconnected components, suggesting hidden links, so we restrict our analysis to cascades occurring entirely within the LSCC. The resulting graph contains 220,974 nodes, 16,772,693 edges, and 32,313 cascades. Initial opinions are set using the polarizing vectors described in Section III-D. However, their computation requires computing the eigensystem of an $n\times n$ dense matrix, which becomes impractical due to the extensive size of the network, even when employing optimized algorithms. To tackle this challenge, we derived a subgraph known as Nepal top cascades, which includes nodes from the 500 longest cascading events. The rationale behind focusing on the longest cascades is that they inherently present a greater challenge for our model, as they involve more complex and prolonged interactions among nodes and increase the likelihood that cascades can overlap (thus, violating assumption (A3)). Specifically, we ensured that all nodes participating in these cascades are included in the subgraph using a heuristic for the directed Steiner Tree problem [54]. Table II summarizes the key properties of this subgraph, which encompasses over 6,000 cascades. To assess the robustness of our results with respect to the selection of the top-cascade subgraph, we replicated the full experimental analysis on an additional sampled subgraph; the corresponding results, which are qualitatively consistent with those reported here, are presented in the SI.
TABLE II: Summary of the real-world Nepal dataset subgraphs used for validation
@@
Our goal is to evaluate how well this real dynamic is captured by the proposed FJC model (we also include the classical FJ model for reference). For the FJC model, we need to estimate the resharing probability $\theta_{i}$ for each node. Inspired by weighted cascade models [33, 49], we define $\theta_{i}$ as the ratio between the number of posts reshared by node $i$ and the number of posts it has seen from its followees. Table III shows the distribution of $\theta_{i}$, which is generally low, indicating that most nodes rarely reshare posts. Please note that these values are consistent with the retweet probabilities in [45], remaining below 0.02. Unlike the real dynamics, the FJC model processes cascades in a tree manner and sequentially. Specifically, posts are propagated from the sources to the network layer-by-layer, in accordance with assumption (A1) and each cascade starts only after the previous one ends, in accordance with assumption (A3).
TABLE III: Statistics of followers, followed and $\theta$ in the Nepal subgraph
@@ V-B2 Opinion evolution of real opinion dynamics vs FJC and FJ
Due to the large size of the networks, we cannot visualize the polarizing vectors, since the plots would be unreadable. Instead, our main objective with these datasets is to evaluate the robustness of the FJC model against the real opinion dynamics. We do so by comparing the distributions of opinion shifts—i.e., the differences between final and initial opinions—for the real evolution (labeled Real) and the predictions of FJC and FJ. As shown in Figure 3, FJC closely mirrors the real dynamics observed in the Nepal top cascades dataset, with the FJC and Real curves largely overlapping, outperforming the standard FJ model. Please note that long cascades are more likely to violate assumption (A3), reducing FJC’s accuracy. Nevertheless, FJC remains surprisingly reliable even in this more challenging scenario.